# Projections of standardised energy indices in future climate scenarios

### v1

@@ 4.2 Energy systems adaptation to mitigate EDs
The previous section illustrated that solar power production droughts are projected to become more common, persistent, and intense as a result of climate change, particularly in central and northern Europe. In this section, we explore two possible ways to mitigate the changes in EDs: increasing the installed capacity, and varying the contribution of solar and wind ($a$) to the hybrid solar-wind generator.
@@
Increasing the installed wind and solar capacities allows us to study how much additional installed capacity is required for EDs to not become more common or intense in the future. This will be important for policymakers when deciding how toconfigure energy systems in the future. We multiply the capacity factors in the future climate by a factor of 1.1 (10% increase), 1.25 (25% increase), and 1.5 (50% increase). The present-day time series remains unaltered, since we want to study how changes in the energy drought system can affect the resulting drought characteristics relative to the current energy system.
@@
Figure 6 shows the changes in the frequency of future wind droughts for the different hypothetical increases of installed capacities. Results are shown as a percentage decrease relative to when the installed capacity is unaltered. Unsurprisingly, increasing the installed capacity severely decreases the frequency of future energy droughts, particularly in southern Europe during summer. Similar changes are found for SPV, with a smaller reduction in winter and autumn in northern Europe (see supplementary material, Figure. S9).
Figure 6: Projected changes in energy drought days for WON generation relative to the unaltered future capacity factor. CMIP5 multimodel mean for RCP8.5 scenario (2048-2098). Assuming (a,b,c,d) 10%, (e,f,g,h) 25% and (i,j,k,l) 50% increase of future capacity factor. Hatches denote regions of ensemble agreement on the sign of bias; i.e., more than 6 of the ensemble members indicate a bias of the same sign.
@@
Figure 7: Projected changes in moderate energy drought days for Hybrid generation. CMIP5 multimodel mean for RCP8.5 scenario (2048-2098). Future shared fraction ($a$) of (a,b,c,d) 20%, (e,f,g,h) 50% and (i,j,k,l) 80%. Hatches denote regions of ensemble agreement on the sign of bias; i.e., more than 6 of the ensemble members indicate a bias of the same sign.
5 Conclusions
In this study, we analyse projected changes in the frequency, duration, and intensity of wind and solar production droughts using standardised energy indices, and investigate how potentially impactful changes can be mitigated by increasing the installed capacity or adapting the mix of renewable energy sources in future energy systems. Our results support previous conclusions regarding the frequency and duration of energy droughts under future climate projections, but by employing the SREPI, we can additionally compare changes to the magnitude and intensity of energy droughts in different regions.
Using CMIP5 climate projections in Europe, we find that solar production will be more heavily affected by changes to the climate than wind production, particularly at high latitudes. This results in solar production droughts in Scandinavia that occur more frequently, persist for longer, and are more intense. We illustrate that this coincides with projected increases in total cloud cover over this region, which aligns with previous research on the behaviour of storm tracks and extratropical cyclones in Europe. In contrast, the increased solar radiation caused by climate change will result in greater solar power production in southern European countries, decreasing the impacts associated with energy droughts.
This motivates future studies on the interaction between weather patterns and future energy droughts. Extratropical cyclones are projected to shift eastward in the North Atlantic basins, resulting in increased cloud clover over northern Europe (Zappa et al., 2013; Dolores-Tesillos et al., 2022; Priestley and Catto, 2022), but energy droughts are also affected by atmospheric blocking, which does not exhibit clear patterns in future climate scenarios (Woollings et al., 2018; Davini and D’Andrea, 2020). Future work into the effects of atmospheric blocking on energy droughts in future climate projections would therefore help to provide additional meteorological reasoning for the results seen herein. In particular, it would help to attribute changes in wind power production, which is affected by multiple convoluted factors. This could additionally be studied in the context of compound drought events (co-occurrence of wind and solar droughts), which have the potential to generate particularly large impacts on energy systems.
@@ 5 Conclusions
We hope that this analysis will be useful for policymakers to understand where and how installed capacities should be increased to counteract the effects of climate change. However, for such studies to be more directly relevant for policymaking, we argue that more publicly available data is required. The study herein does not consider off-shore data, despite the large number of wind and solar power plants based off-shore, and only considers wind and solar power, despite other sources of renewable energy contributing to national energy mixes, most notably hydropower. We also make the simplifying assumption that a larger wind speed results in a larger wind power production, neglecting the fact that wind turbines often cannot operate under high wind speeds; more data is required to incorporate this into the analysis.
@@
Future work would also benefit from having up-to-date installed wind and solar capacities available, at either a regional or national level. This would allow us to analyse energy production droughts in systems with both wind and solar production. The hybrid generator defined in Equation 1 provides a simple approach to simulate this, but the assumption that wind and solar each contribute 50% to the renewable energy mix (in all regions) is somewhat arbitrary, and is not always representative of the operational set up in the regions considered. However, one value must be chosen in the absence of regional installed capacity data. We therefore advocate further work to make such data available, which we believe would benefit the practical applicability of research results to policymaking.

### v2

@@ 4.3 Energy system adaptation
The previous section illustrated that solar power production droughts are projected to become more common, persistent, and intense as a result of climate change, particularly in central and northern Europe. In this section, we explore two possible ways to mitigate the changes in energy droughts: increasing the installed capacity, and varying the contribution of solar and wind ($a$) to the hybrid solar-wind generator.
@@
Increasing the installed wind and solar capacities allows us to study how much additional installed capacity is required for energy droughts to not become more common or intense in the future. This will be important for policymakers when deciding how to configure energy systems in the future. We multiply the capacity factors in the future climate by a factor of 1.1 (10% increase), 1.25 (25% increase), and 1.5 (50% increase). The present-day time series remains unaltered, since we want to study how changes in the energy drought system can affect the resulting drought characteristics relative to the current energy system.
@@
Changes in the frequency of future wind droughts for the different hypothetical increases of installed capacities are shown in supplementary material (Figs. S15 and S16). Results are shown as a percentage decrease relative to when the installed capacity is unaltered. Unsurprisingly, increasing the installed capacity severely decreases the frequency of future energy droughts, particularly in southern Europe during summer. Similar changes are found for SPV, with a smaller reduction in winter and autumn in northern Europe (see supplementary material, Figure. S16).
@@
Figure 6: Projected changes in moderate energy drought days for Hybrid generation. CMIP5 multimodel mean for RCP4.5 scenario (2048-2098). Future shared fraction ($a$) of (a,b,c,d) 20%, (e,f,g,h) 50% and (i,j,k,l) 80%. Each column corresponds to a season: a,e,i) winter, b,f,j) spring, c,g,k) summer and d,h,i) autumn. Hatches denote regions of ensemble agreement on the sign of bias; i.e., more than 6 of the ensemble members indicate a bias of the same sign.
5 Summary
In this study, we analyse projected changes in the frequency, duration, and intensity of wind and solar production droughts using standardised energy indices, and investigate how potentially impactful changes can be mitigated by increasing the installed capacity or adapting the mix of renewable energy sources in future energy systems. Our results support previous conclusions regarding the frequency and duration of energy droughts under future climate projections, but by employing the SREPI to define energy droughts, we can additionally compare changes to the magnitude and intensity of energy droughts in different regions.
Using EURO-CORDEX climate simulations (based on CMIP5), we find that solar production will be more heavily affected by changes to the climate than wind production, particularly at high latitudes. This results in solar production droughts in Scandinavia that occur more frequently, persist for longer, and are more intense. We illustrate that this coincides with a projected increase in total cloud cover over this region, which aligns with previous research on the behaviour of storm tracks and extratropical cyclones in Europe. In contrast, the increased solar radiation caused by climate change will result in greater solar power production in southern European countries, decreasing the impacts associated with energy droughts.
This motivates future studies on the interaction between weather patterns and future energy droughts. Extratropical cyclones are projected to shift eastward in the North Atlantic basins, resulting in increased cloud clover over northern Europe (Zappa et al., 2013; Dolores-Tesillos et al., 2022; Priestley and Catto, 2022), but energy droughts are also affected by atmospheric blocking, which does not exhibit clear patterns in future climate scenarios (Woollings et al., 2018; Davini and D’Andrea, 2020). Future work into the effects of atmospheric blocking on energy droughts in future climate projections would therefore help to provide additional meteorological reasoning for the results seen herein.In particular, it would help to attribute changes in wind power production, which is affected by multiple convoluted factors. Moreover, additional efforts should focus on quantifying the impacts of atmospheric blocking on energy demand under climate change, which is greatly influenced by weather patterns (Otero et al., 2022). This could additionally be studied in the context of compound drought events (co-occurrence of wind and solar droughts), which have the potential to generate particularly large impacts on energy systems.
@@ 5 Summary
6 Discussion: The need for more energy data
The goal of this analysis is to provide information that may be useful for policymakers to understand what action, if any, is required to counteract the effects of climate change on renewable energy systems. However, for such studies to be more directly relevant for policymaking, we argue that more publicly available energy data is needed. We therefore end this paper with a discussion of the limitations of this study that arise due to a lack of data, and advocate further work to make such data available.
@@ 6 Discussion: The need for more energy data
The study herein does not consider off-shore data, despite the large number of wind and solar power plants based off-shore, and only considers wind and solar power, despite other sources of renewable energy contributing to national energy mixes, most notably hydropower. We have limited information regarding the physics and management of power systems, including the efficiency of renewable plants when harvesting wind and solar energy, the plants’ operational hours, and their ability to store energy. Similarly, we make the simplifying assumption that higher wind speeds lead to increased wind power production, disregarding that wind turbines often cannot operate under high wind speeds. Additional data is required to incorporate these aspects into the analysis, and it would therefore be incredibly beneficial to have a unified data set that contains all of this information.
Furthermore, our results are specific for each region considered in the study. That is, we assume that the risks of a power outage are determined uniquely by the energy production at that region. Such a framework allows us to analyse local characteristics of the energy system, acknowledging that the climate, and hence renewable energy production, will evolve differently at a sub-country level. However, energy systems are typically operated at a national level, with energy sharing between countries also possible. Having a reasonable representation of how renewable energy resources are shared between regions and countries would permit a more realistic assessment of the risks of energy droughts. Future studies could then investigate how energy sharing strategies could be adapted to reduce the risks of energy droughts.
We restrict attention throughout to renewable energy production droughts based on the SREPI. However, van der Wiel et al. (2019) additionally study energy supply droughts that depend on the residual load (demand minus production), which can similarly be defined in terms of the Standardised Residual Load Index (SRLI; Allen and Otero, 2023b). Since energy demand is itself highly weather dependent, it would be informative to study projections of the SRLI in future climate scenarios, which would allow us to analyse how energy supply droughts are expected to evolve in the future. However, such an analysis is currently not possible, since we do not have access to up-to-date installed capacity data in each region of interest, and therefore cannot calculate the total renewable energy production.
@@
Future work would therefore greatly benefit from having up-to-date installed wind and solar capacities available, at either a regional or national level. The hybrid generator defined in Equation 1 provides a simple approach to simulate the total renewable energy production (Kapica et al., 2024), but the assumption that wind and solar each contribute 50% to the renewable energy mix (in all regions) is somewhat arbitrary, and is generally not representative of the operational set up in the regions considered. However, this provides a necessary and useful baseline in the absence of specific regional capacity data.
Finally, an analysis of energy supply droughts would also require an accurate estimate of energy demand in the future. Energy demand is typically estimated from weather data via the number of hot and dry days, though this additionally introduces uncertainty that should be accounted for. In particular, such an approach neglects other important predictors of energy demand, such as the prevailing socio-economic conditions. However, designing a more complex model for energy demand is hindered by the lack of prolonged, historical time series of regional energy demand. Similarly, there are also uncertainties in the future climate simulations, as well as the corresponding energy indicators from the Copernicus Climate Data Store (2021). While these have been bias-corrected, some biases in the EURO-CORDEX data may remain; Bartók et al. (2017) find that solar irradiance is generally over-estimated, for example.
We are very grateful to Hannah Bloomfield and Jacek Kapica for useful comments during the preparation of this work. EDT acknowledgments the nextGEMS project.